OpenStudy (aleah54):

Choose the correct simplification of the expression (3x)4.

OpenStudy (misty1212):

hI!!

OpenStudy (misty1212):

is it \((3x)^4\)??

OpenStudy (misty1212):

if so you can rewrite it as \[(3x)^4=3^4x^4=81x^4\] maybe that is your answer

OpenStudy (aleah54):

how did you get that answer

OpenStudy (aleah54):

@misty1212

alones (alones):

Aleah was it \(\ (3x)(4)\)? or it was \(\ (3x)^4\)

alones (alones):

Since if it's \(\ (3x)(4\)) than the right answer would be \(\ 12x\) and if it's a second option than misty is right.

OpenStudy (aleah54):

the answer was 81x^4

jhonyy9 (jhonyy9):

Aleah please clarify how is exactly your exercise bc. this 81x^4 is the right answer just than there are (3x)^4 wrote

jhonyy9 (jhonyy9):

bc. this result 3^4 = 81 and than x^4 = togheter 81x^4 hope helped understanding